OpenStudy (shaleiah):

A race car accelerates at a rate of 15.6 m/s^2. If it starts from rest, how much time till it is traveling at 325 km/h?

OpenStudy (welshfella):

use one of the equations for motion under constant acceleration;- v = u + at u = 0 , v = 325, a = 15.6 , t = ? solve for t.

OpenStudy (welshfella):

plug in those values and solve for t:- 325 = 0 + 15.6t

OpenStudy (shaleiah):

340.6

OpenStudy (welshfella):

no 15.6t = 325 divide both sides by 15.6

OpenStudy (shaleiah):

20.8333

OpenStudy (welshfella):

you need to find the value of t